# Mechanical Properties of the Proton from a Deformed AdS Holographic Model

### v1

@@
Abstract
@@ I Introduction
The proton, one of the building blocks of nuclear matter and, consequently, ordinary stuff, has been known for more than one century; however, some of its properties remain at least partially unknown. This is due mainly to its internal structure with non-trivial contributions from its constituents, the quarks and gluons. One way to address this endeavor is to look at the Energy Momentum Tensor’s (EMT) gravitational form factors (GFFs), which encode information regarding the distribution of its mass or momentum, spin, shear, and pressure. This question naturally expands to nucleons and other hadrons Burkert:2018bqq; Polyakov:2018zvc; Deur:2018roz; Lorce:2018egm; Burkert:2023wzr. In principle, GFFs can be probed by the gravitational interaction of quarks and gluons with the graviton, but this constitutes a challenging experimental effort because of the weakness of the gravitational interaction at the quantum level. This was proposed long ago by Gell-Mann Gell-Mann:1962yej. Recently, experimental extractions of GFFs have been performed indirectly, for instance, accessing them through deep virtual Compton scattering data Burkert:2018bqq; Burkert:2021ith. The future Electron Ion Collider Accardi:2012qut looks to further extend such experimental investigations using scattering processes to extract GFFs and to investigate the nucleon structure and its properties. On the theory side, recent developments in lattice QCD, such as Alexandrou:2018xnp; Shanahan:2018pib; Shanahan:2018nnv; Hackett:2023rif, explored proton GFFs and some of its mechanical properties, such as pressure and shear distributions.
@@
As non-perturbative aspects of QCD related to its strong coupling regime, such properties of the nucleon are also well suited for holographic models based on the gauge/gravity duality Maldacena:1997re; Gubser:1998bc; Witten:1998qj, which relates a classical weakly coupled gravity theory on a curved five dimensional anti-de Sitter space to a dual strongly coupled gauge theory on flat Minkowski space. In this context, the works of Abidin:2008ku; Abidin:2008hn; Brodsky:2008pf; Abidin:2008sb; Abidin:2009hr; Anderson:2014jia; Chakrabarti:2015lba; Mondal:2015fok; Mondal:2017lph; Kumar:2017dbf; Watanabe:2018owy; Mamo:2019mka; Aliev:2020aih; Chakrabarti:2020kdc; Mamo:2021krl; Karapetyan:2021crv; Liu:2022out; Allahverdiyeva:2023fhn; Mamedov:2024tth; Deng:2025fpq; Sain:2025kup paved the way for studies of nucleon and hadronic GFFs and some of their mechanical properties such as those mentioned above using holographic models for QCD, also known as AdS/QCD models. Among these models, we mention the hard wall Polchinski:2001tt; Boschi-Filho:2002xih; Boschi-Filho:2002wdj; deTeramond:2005su; Erlich:2005qh; Boschi-Filho:2005xct, the soft wall Karch:2006pv; Brodsky:2006uqa; Erdmenger:2007cm; Brodsky:2014yha, which are classified as bottom-up models, as well as some top-down models such as the D4D8D4-D8 Sakai-Sugimoto Sakai:2004cn; Sakai:2005yt; Hata:2007mb. The role of confinement in the soft wall model inspired Andreev:2006ct another bottom-up AdS/QCD model known as the deformed AdS FolcoCapossoli:2019imm; FolcoCapossoli:2020pks; MartinContreras:2021yfz; Contreras:2021epz. A comparison remarking similarities and differences of soft wall and deformed AdS models appeared recently in Nascimento:2023dzx.
@@
In this work, we explore the proton gravitational form factors by computing the eigen values of the nucleon EMT in a deformed AdS background holographic model FolcoCapossoli:2019imm; FolcoCapossoli:2020pks; MartinContreras:2021yfz; Contreras:2021epz. One of them, the $D$ term, allows one to explore mechanical properties of the system such as pressure and shear distributions, as well as internal forces.
@@
This paper is organized as follows. In Section II, we describe the holographic model used in this work and how to describe the nucleon as a Dirac field on this background. Section III is devoted to the computation of the gravitational form factors of the nucleon, and the comparison of our results with some lattice QCD data, as well as with the holographic results of Mamo:2019mka. From these, in Section IV, we explore the pressure, shear, and energy distributions in the nucleon using a dipole approximation to our numerical results, which also allows us to compute the mechanical radius associated with these distributions. Using these results, in Section V, we investigate the internal forces in the proton, such as normal and tangential forces inside of it. We draw our conclusions and future perspectives in Section VI.
II Holographic Model
@@
In the deformed AdS model, we consider a five-dimensional anti-de Sitter background deformed by a warp factor $\mathcal{A}(z)$ as follows Contreras:2021epz; Nascimento:2023dzx
@@
$\mathcal{A}(z)=-\log z+\frac{1}{2}k\,z^{2},$ (2)
where $\log z$ refers to the asymptotic background $z\to 0$ $\text{AdS}_{5}$, with radius of curvature $R=1$, which, in the bulk, is deformed by a dilaton-like factor of $\frac{1}{2}k\,z^{2}$, with $k$ a constant having a mass squared dimension. The extra term in the deformed AdS background is responsible for breaking the conformal invariance of the boundary quantum field theory. It was originally chosen to reproduce confining quark-antiquark potentials Andreev:2006ct.
II.1 Proton in the deformed AdS background
Fermions, such as the proton, are described by a probe Dirac field $\Psi$ in five dimensions Contreras:2021epz; Nascimento:2023dzx
@@ II.1 Proton in the deformed AdS background
with $A^{\prime}(z)$ denoting the derivative with respect to the holographic coordinate $z$. We can decompose $\Psi$ into left e right chiral components
$\Psi(x^{\mu},z)=\Psi_{{}_{L}}(x^{\mu},z)+\Psi_{{}_{R}}(x^{\mu},z),,$ (9)
@@
with $M_{n}$ the mass of the $n$th excited state. By decoupling these equations and using the following transformations,
@@
In order to model gravitational form factors, we need to solve (14) for its ground state. With $\mathcal{A}(z)$ given by (2), the resulting fermionic solution is obtained by numerically solving the Schrödinger-like equation. The goal is to make the ground state mass $M_{0}$ match the proton mass of $M_{p}=0.938\,\text{GeV}^{2}$. We seek for free parameters to achieve this goal. The constant $k$ present in the background warp factor $\mathcal{A}(z)$ can be fine tuned along with the bulk mass $m_{5}$, given by
$m_{5}=\Delta-2,$ (16)
where $\Delta$ is the fermionic conformal dimension $\Delta=\frac{3}{2}$ Contreras:2021epz; Nascimento:2023dzx. With the resulting value of $m_{5}$ we are unable to reproduce the mass of the proton for any chosen value of $k$. For this purpose, we use an anomalous dimension $\gamma$ such that
$m_{5}=\Delta+\gamma-2.$ (17)
In this way, we can fine tune both parameters, $k$ and $\gamma$, to not only match $M_{0}$ with the proton mass, but to adjust the predictions of this model for the gravitational form factors of the proton with the corresponding results in the lattice QCD literature, and compare them with previous holographic results. As we shall see in the next section, the values that best achieve this goal are found to be $k=(0.343)^{2}$$\text{GeV}^{2}$ and $\gamma=1.745$, as illustrated by the masses of the $\psi_{L}$ e $\psi_{R}$ solutions in Table 1.
@@
$M_{0}$ $0.937801\,\text{GeV}^{2}$ $0.937135\,\text{GeV}^{2}$
Table 1: Masses of the chiral solutions $\psi_{L}$ and $\psi_{R}$ for $k=(0.343)^{2}\,\text{GeV}^{2}$ and $\gamma=1.745$.
@@ III Gravitational Form Factors
The mechanical properties of matter are related to the way in which it interacts with gravitational fields Burkert:2023wzr. The response of gravity to the presence of matter is translated into the energy-momentum tensor (EMT) such that matter mechanical properties can be extracted from its matrix elements between single particle states in terms of form factors. The general decomposition of QCD EMT in terms of gravitational form factors is given by Polyakov:2018zvc; Mamo:2019mka; Mamo:2021krl
@@
with $a^{(\mu}b^{\nu)}=\frac{1}{2}(a^{\mu}\,b\nu+a^{\nu}\,b^{\mu})$, $q^{2}=(p_{2}-p_{1})^{2}$ and $\bar{u}u=2\,m_{N}$, with $m_{N}$ the nucleon mass. The scalar coefficients $A$, $B$, and $C$ are the gravitational form factors, which encode information on momentum distribution, spin and angular momentum, as well as internal mechanical forces distributions inside the nucleon, respectively.
@@
$T^{\mu\nu}(x)=\frac{2}{\sqrt{-g}}\frac{\delta\,S_{\text{grav}}}{\delta\,g^{\mu\nu}(x)},$ (19)
@@
$g_{MN}(z)\to g_{MN}(z)+h_{MN}(z),$ (20)
with line element given by (1), and warp factor $\mathcal{A}(z)=-\log z+k\,z^{2}/2$. The form factors are obtained by coupling the metric fluctuation $h_{MN}$ to the Dirac fermions $\psi_{L/R}$ present in the bulk EMT. To reproduce its tensorial structure in (18), we decompose the metric fluctuation into a transverse-traceless part $h(q,z)$, and a longitudinal trace full part $H(q,z)$
$h_{MN}(q,z)=\epsilon^{TT}_{MN}\,h(q,z)+q_{M}\,q_{N}\,H(q,z).$ (21)
In the axial gauge $h_{Mz}=h_{z}z=0$, the corresponding graviton equation of motion obtained from the linearized Einstein’s equations takes the form of
$\frac{1}{\sqrt{-g}}\,\partial_{M}(\sqrt{-g}\,g^{MN}\partial_{N}\,h_{\mu\nu})=0.$ (22)
By numerically solving this last equation for each component of $h_{\mu\nu}$, one can obtain the form factors using (3) as the bulk gravitational action in (19). As a consequence of this choice, one can show that there will be no coupling between metric fluctuation and a spin term $\sigma^{\mu}$ in the deformed AdS model, resulting in a vanishing $B$ form factor.

### v2

@@
Mechanical Properties of the Proton from a Deformed AdS Holographic Model
Ayrton Nascimento1 ayrton@pos.if.ufrj.br    Henrique Boschi-Filho1 boschi@if.ufrj.br 1Instituto de Física, Universidade Federal do Rio de Janeiro, 21.941-909, Rio de Janeiro, RJ, Brazil
Resumo
@@ I Introduction
The proton, one of the building blocks of nuclear matter and, consequently, ordinary stuff, has been known for more than one century. However, some of its properties remain at least partially unknown. This is due mainly to its internal structure with non-trivial contributions from its constituents, the quarks and gluons. One way to address this endeavor is to look at the Energy Momentum Tensor’s (EMT) gravitational form factors (GFFs), which encode information regarding the distribution of its mass or momentum, spin, shear, and pressure. This question naturally expands to nucleons and other hadrons Burkert et al. (2018); Polyakov and Schweitzer (2018); Deur et al. (2019); Lorcé et al. (2019); Burkert et al. (2023). In principle, GFFs can be probed by the gravitational interaction of quarks and gluons with the graviton, but this constitutes a challenging experimental effort because of the weakness of the gravitational interaction at the quantum level. This was proposed long ago by Gell-Mann Gell-Mann (1962). Recently, experimental extractions of GFFs have been performed indirectly, for instance, accessing them through deep virtual Compton scattering data Burkert et al. (2018, 2021). The future Electron Ion Collider Accardi and others (2016) looks to further extend such experimental investigations using scattering processes to extract GFFs and to investigate the nucleon structure and its properties. On the theory side, recent developments in lattice QCD, such as Alexandrou et al. (2018); Shanahan and Detmold (2019a, b); Hackett et al. (2024), explored proton GFFs and some of its mechanical properties, such as pressure and shear distributions.
@@
As non-perturbative aspects of QCD related to its strong coupling regime, such properties of the nucleon are also well suited for holographic models based on the gauge/gravity duality Maldacena (1998); Gubser et al. (1998); Witten (1998), which relates a classical weakly coupled gravity theory on a curved five dimensional anti-de Sitter space to a dual strongly coupled gauge theory on flat Minkowski space. In this context, the works of Abidin and Carlson (2008b, a); Brodsky and de Teramond (2008); Abidin and Carlson (2008c, 2009); Anderson et al. (2014); Chakrabarti et al. (2015); Mondal (2016); Mondal et al. (2017); Kumar et al. (2017); Watanabe and Huang (2019); Mamo and Zahed (2020); Aliev et al. (2021); Chakrabarti et al. (2020); Mamo and Zahed (2021); Karapetyan (2022); Liu et al. (2022); Allahverdiyeva and Mamedov (2023); Mamedov et al. (2025); Deng and Hou (2025); Sain et al. (2025) paved the way for studies of nucleon and hadronic GFFs and some of their mechanical properties such as those mentioned above using holographic models for QCD, also known as AdS/QCD models. Among these models, we mention the hard wall Polchinski and Strassler (2002); Boschi-Filho and Braga (2003, 2004); de Teramond and Brodsky (2005); Erlich et al. (2005); Boschi-Filho et al. (2006), the soft wall Karch et al. (2006); Brodsky and de Teramond (2006); Erdmenger et al. (2008); Brodsky et al. (2015) (see also Abidin and Carlson (2008a, c, b, 2009); Mamo and Zahed (2020, 2021); Braga and Vega (2012)), which are classified as bottom-up models, as well as some top-down models such as the D4D8D4-D8 Sakai-Sugimoto Sakai and Sugimoto (2005a, b); Hata et al. (2007). The role of confinement in the soft wall model inspired Andreev and Zakharov (2006) another bottom-up AdS/QCD model known as the deformed AdS Folco Capossoli et al. (2020a, b); Martin Contreras et al. (2022); Contreras et al. (2021). A comparison remarking similarities and differences of the soft-wall and deformed AdS models appeared recently in Nascimento and Boschi-Filho (2023). The soft-wall model considers the pure AdS space and the introduction of the exponential of a dilaton field in the action Karch et al. (2006). In the case of fermions it is also necessary to introduce a potential in the mass term to obtain discrete mass spectra Abidin and Carlson (2008a, b). On the other hand, the deformed AdS model considers an exponential with the dilaton field in the metric, which is then only asymptotically AdS and there is no extra dilaton in the action Folco Capossoli et al. (2020a, b). In fact, in the soft-wall model, the dilaton-like term present in the fields’ action appears through an exponential that factors out of the equations of motion, making no effect on the fermion field dynamics. It also does not affect the fields’ mass spectra. Introducing a $z$ dependent mass term in the action modifies the fields’ dynamics and, together with the dilaton, leads to a discrete spectrum of eigenvalues. Regarding the deformed AdS model, in which the dilaton-like term goes into the line element of the asymptotic AdS background, there is no need for an additional mass term proportional to the $z$ coordinate in the Dirac action, leaving it more natural in the bulk geometry. As a consequence, a discrete spectrum naturally emerges. In addition, Schrödinger-like equations of motion coming from a bulk Dirac action present numerical solutions, as opposed to the analytical solutions present in the soft-wall model in Abidin and Carlson (2008a); Mamo and Zahed (2020). In fact, as we have shown in Nascimento and Boschi-Filho (2023) in the case of fermions, in the deformed AdS model, in a quadratic approximation, the equations of motion can be cast in a way that look like those of the soft-wall model, leading to similar analytical solutions, but even in this case their spectra differ by the ground state, which shows the models, in this approximation, may look similar to one another, but are fundamentally different.
@@
In this work, we explore the proton gravitational form factors by computing the eigenvalues of the nucleon EMT in a deformed AdS background holographic model Folco Capossoli et al. (2020a, b); Martin Contreras et al. (2022); Contreras et al. (2021). One of them, the $D$ term, allows one to explore mechanical properties of the system such as pressure and shear distributions, as well as internal forces.
@@
This paper is organized as follows. In Section II, we discuss the deformed AdS holographic model used in this work and describe the nucleon as a Dirac field in the background spacetime of this model. Section III is devoted to the computation of the gravitational form factors of the nucleon, and the comparison of our results with some lattice QCD data, as well as with other holographic results. From these, in Section IV, we explore the pressure, shear, and energy distributions in the nucleon using a dipole approximation to our numerical results, which also allows us to compute the mechanical radius associated with these distributions. Using these results, in Section V, we investigate the internal forces in the proton, such as normal and tangential forces inside of it. We draw our conclusions and future perspectives in Section VI. We also include an Appendix A where we briefly review the usual soft wall model for fermions, pinpointing the differences with the deformed AdS model discussed in Section II.
II Deformed AdS model
@@
In the deformed AdS model, we consider a five-dimensional anti-de Sitter background deformed by a warp factor $\mathcal{A}(z)$ as follows Contreras et al. (2021); Nascimento and Boschi-Filho (2023)
@@
$\mathcal{A}(z)=-\log\left(\frac{z}{R}\right)+\frac{1}{2}k\,z^{2},$ (2)
where $\log\left(\frac{z}{R}\right)$, with $R$ the radius of curvature of the asymptotic $\text{AdS}_{5}$ background, recovered within the limit $z\to 0$. The dilaton-like factor of $\frac{1}{2}k\,z^{2}$, with $k$ a constant having a mass squared dimension, deforms the AdS background breaking the conformal invariance of the boundary quantum field theory. It was originally chosen to reproduce confining quark-antiquark potentials Andreev and Zakharov (2006).
Fermions, such as the proton, are described by a probe Dirac field $\Psi$ in this model in five dimensions Contreras et al. (2021); Nascimento and Boschi-Filho (2023)
@@
with $\mathcal{A}^{\prime}(z)$ representing the derivative of $\mathcal{A}(z)$ with respect to the holographic coordinate $z$. Note that, according to its definition, Eq. (2), $\mathcal{A}(z)$ is dimensionless, and so $\mathcal{A}^{\prime}(z)$ has dimension of mass, as well as $m_{5}$, introduced in the fermionic action, Eq. (3) (see also the discussion below Eq. (17) where we set the value of $m_{5}$ in GeV, together with other parameters in the model to fit the proton mass.
In order to solve Eq. (8), we can decompose $\Psi$ into its left and right chiral components
$\Psi(x^{\mu},z)=\Psi_{{}_{L}}(x^{\mu},z)+\Psi_{{}_{R}}(x^{\mu},z),$ (9)
@@
with $M_{n}$ the mass of the $n$th excited hadronic state in four dimensions. The one corresponding to the ground state $M_{0}$ is to be identified with the proton’s mass. By decoupling these equations and using the following transformations,
@@
In order to model gravitational form factors, we need to solve (14) for its ground state. With $\mathcal{A}(z)$ given by (2), the resulting fermionic solution is obtained by numerically solving the Schrödinger-like equation. The goal is to make the ground state mass $M_{0}$ match the proton mass of $M_{p}=0.938\,{\text{GeV}}$. We seek for free parameters to achieve this goal. The constant $k$ present in the background warp factor $\mathcal{A}(z)$ could be chosen so that the match to $M_{p}$ is achieved. But this is not accomplished. Another free parameter is needed for this match. This is generally resolved by looking at $m_{5}$ entering (14) through (15). In particular, according to the holographic dictionary, in the pure AdS spacetime, $m_{5}^{\text{AdS}}$ is related to the canonical scaling conformal dimension $\Delta_{\text{can}}$ of a boundary operator $\mathcal{O}$ through
$R\,m_{5}^{\text{AdS}}=\Delta_{\text{can}}-2,$ (16)
where $\Delta_{\text{can}}$ is the fermionic canonical conformal dimension $\Delta=\frac{3}{2}$. With the resulting value of $m_{5}$, we are unable to reproduce the mass of the proton for any chosen value of $k$ when using it in solving Eq. (14). In fact, in Georgi and Politzer (1974); Gross and Wilczek (1974); Gross et al. (1977); Jaroszewicz (1982) it was shown that the canonical dimension of an operator $\mathcal{O}$ should be modified by introducing an anomalous dimension $\gamma$, implying an effective scaling dimension $\Delta_{\text{eff}}=\Delta_{\text{can}}+\gamma-2$. Inspired by this, we consider an anomalous dimension $\gamma$ in $m_{5}$ such that
$R\,m_{5}=\Delta_{\text{can}}+\gamma-2.$ (17)
In this way, we can fine tune both parameters, $k$ and $\gamma$, to not only match $M_{0}$ with the proton mass, but to adjust the predictions of this model for the gravitational form factors of the proton with the corresponding results in the lattice QCD literature, and compare them with previous holographic results. To solve (14), we choose $R=1\,\text{GeV}^{-1}$ and set $z_{0}=1\times 10^{-5}\,\text{GeV}^{-1}$ and $z_{\infty}=7\,\text{GeV}^{-1}$ as the domain to which constrain the solutions of (14). The values of the free parameters that best achieve the desired match between $M_{0}$ and $M_{p}$ are found to be $k=(0.343)^{2}$$\text{GeV}^{2}$, $\gamma=1.745$ and $m_{5}=1.245\,\text{GeV}$, as illustrated by the masses $M_{0}$ of the $\psi_{L}$ and $\psi_{R}$ solutions in Table 1.
@@
$M_{0}$ $0.937801\,{\text{GeV}}$ $0.937135\,{\text{GeV}}$
Tabela 1: Masses of the chiral solutions $\psi_{L}$ and $\psi_{R}$ for $k=(0.343)^{2}\,\text{GeV}^{2}$, $\gamma=1.745$ and $m_{5}=1.245\,\text{GeV}$.
The corresponding chiral potentials for (15) present a confining behavior, as shown in Fig. 1.
Figura 1: Confining chiral potentials $V_{L/R}(z)$ which come from (15) with the free parameters given by the particular choice $k=(0.343)^{2}$$\text{GeV}^{2}$ and $\gamma=1.745$.
The resulting ground state solutions $\psi_{L}$ and $\psi_{R}$ corresponding to the eigenvalues in Table 1 are depicted in Fig. 2. In what follows, we keep all the parameters discussed above fixed.
Figura 2: Ground state solutions $\psi_{L}(z)$ and $\psi_{R}(z)$ corresponding to the chiral potentials $V_{L/R}(z)$. We associate these solutions with the proton wave function $\psi_{p}(z)$ in its ground state in the present holographic model.
@@ III Gravitational Form Factors
The mechanical properties of matter are related to the way in which it interacts with gravitational fields Burkert et al. (2023). The response of gravity to the presence of matter is translated into the energy-momentum tensor (EMT) such that matter mechanical properties can be extracted from its matrix elements between single particle states in terms of form factors. The general decomposition of QCD EMT in terms of gravitational form factors is given by Polyakov and Schweitzer (2018); Mamo and Zahed (2020, 2021)
@@
with kinematic variables $P=\frac{1}{2}(p_{2}+p_{1})$ and $q=(p_{2}-p_{1})$. We also have $a^{(\mu}b^{\nu)}=\frac{1}{2}(a^{\mu}\,b\nu+a^{\nu}\,b^{\mu})$, and $\bar{u}u=2\,m_{N}$, with $m_{N}$ the nucleon mass. We consider the usual spin operator $\sigma^{\mu\nu}=\frac{i}{2}\{\gamma^{\mu},\gamma^{\nu}\}$. The scalar coefficients $A$, $B$, and $C$ are the gravitational form factors, which encode information on momentum distribution, spin and angular momentum, as well as internal mechanical forces distributions inside the nucleon, respectively.
@@
$T^{\mu\nu}(x)=\frac{2}{\sqrt{-g}}\frac{\delta\,S}{\delta\,g^{\mu\nu}(x)},$ (19)
with
$S=S_{\text{G}}+S_{\psi},$ (20)
where
$S_{\text{G}}$ $=\frac{1}{2\,\kappa_{5}^{2}}\int dz\,d^{4}x\,\sqrt{-g}\left(\mathcal{R}-2\,\Lambda\right),$ (21)
[…]
$\mathcal{R}$ is the Ricci scalar, $\kappa_{5}^{2}\equiv 8\pi\,G_{5}$, with $G_{5}$ Newton’s constant in five dimensions, and $\Lambda$ a cosmological constant.
@@
$g_{mn}(x,z)\to g_{mn}(x,z)+h_{mn}(x,z),$ (23)
with line element given by (1), and warp factor $\mathcal{A}(z)=-\log(z/R)+k\,z^{2}/2$. The form factors are obtained by coupling the metric fluctuation $h_{mn}$ to the Dirac fermions $\psi_{L/R}$ present in the bulk EMT. To reproduce its tensorial structure in (18), we decompose the metric fluctuation into a transverse-traceless part $h(q^{2},z)$, and a longitudinal trace full part $H(q^{2},z)$
$h_{mn}(x,z)=\epsilon^{TT}_{mn}\,e^{-i\,q\cdot x}\,h(q^{2},z)+q_{m}\,q_{n}\,e^{-i\,q\cdot x}\,H(q^{2},z),$ (24)
with a plane wave functional form for each of them. In the axial gauge $h_{mz}={h_{zz}}=0$, the corresponding graviton equation of motion obtained from the linearized Einstein’s equations takes the form of
$\frac{1}{\sqrt{-g}}\,\partial_{m}(\sqrt{-g}\,g^{mn}\partial_{n}\,h_{\mu\nu})=0,$ (25)
Since each component in (24) satisfies (25), we shall call them generically $f(q^{2},z)$, which can be both $h(q^{2},z)$ and $H(q^{2},z)$. Also, by taking into account (1) along with (2) in (25), we arrive at the following equation for the graviton modes
$f^{\prime\prime}(q^{2},z)+3\,\mathcal{A}^{\prime}(z)\,f^{\prime}(q²,z)-q^{2}\,f(q^{2},z)=0,$ (26)
where prime denotes derivative with respect to the holographic coordinate, $z$. With the plane wave ansatz, we seek solutions such that $f(q^{2},0)=1$ and $f(q^{2},\infty)=0$. In solving numerically Eqs. (26), we consider setting $q^{2}_{\text{max}}=2\,\text{GeV}^{2}$. This is because we want to compute form factors in the range $0\leq q^{2}\leq 2\,\text{GeV}^{2}$, which corresponds to the kinematical range used in recent lattice QCD calculations, whose results we will compare to ours in the following. In this way, we get the following set of solutions of Eqs. (26) illustrated by Fig. 3.
Figura 3: Graviton solutions $f(q^{2},z)$, which represent both $h(q^{2},z)$ and $H(q^{2},z)$.
With these solutions at hand, we can proceed to compute the the gravitational form factors. One can start by considering (3) as the matter bulk action,
$S_{\psi}=\int dz\,d^{4}x\,e^{5\mathcal{A}(z)}\left[\bar{\Psi}\left(e^{-\mathcal{A}(z)}\gamma^{5}\,\partial_{z}+e^{-\mathcal{A}(z)}\,\gamma^{\mu}\partial_{\mu}+2\,\mathcal{A}^{\prime}(z)\,\gamma^{5}-m_{5}\right)\Psi\right].$ (27)
This enters (19) as $S_{\psi}$, which will give rive to the form factors. As a consequence of this choice, one can show that there will be no coupling between metric fluctuation and a spin term $\sigma^{\mu}$ in the deformed AdS model, resulting in a vanishing $B$ form factor.